OpenStudy (anonymous):

Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (7, 2) (5, 4)

OpenStudy (anonymous):

Start with finding the slope (m) Slope is the change of y over the change of x. The Y values we see are 2 and 4 The X values we see are 7 and 5 So the slope would be \[\frac{ 2-4 }{ 7-5 }\] Solve that and we know the slope is \[\frac{ -2 }{ 2 }\] or -1 So now we have an equation of \[y=-1x+b\] To solve b you simply put in one of your points. Take for example the point of (7,2) So now we have \[2=-1*7+b\] Finish solving for b and then input that into the equation